### v2

# Prototype-Based Disentanglement for Controllable Dysarthric Speech Synthesis

Abstract
Dysarthric speech exhibits high variability and limited labeled data (accurately transcribed recordings), posing major challenges for both automatic speech recognition (ASR) and assistive speech technologies. Current methods depend on synthetic data or reconstructed speech, yet they often fail to separate speaker identity from pathological speech patterns, thereby reducing controllability and robustness.
In this paper, we propose ProtoDisent-TTS, a prototype-based disentanglement TTS framework built on a pre-trained text-to-speech backbone that factorizes speaker timbre and dysarthric articulation within a unified latent space. A pathology prototype codebook (a defined set of speech patterns) provides interpretable and controllable representations of healthy and dysarthric speech, while a dual-classifier objective with a gradient reversal layer ensures speaker identity remains separate from these pathological attributes. Experiments on the TORGO dataset demonstrate that our design can bidirectionally convert healthy and dysarthric speech, resulting in consistent ASR performance gains and robust, speaker-aware speech reconstruction. [^1] Code and audio samples: https://mors20.github.io/ProtoDisent-TTS. The paper has been submitted to Interspeech 2026.
*Correspondence: wangnizhuan1120@gmail.com; wai-ting.siok@polyu.edu.hk
Keywords dysarthric speech synthesis, data augmentation, speech reconstruction, automatic speech recognition (ASR), Text-to-Speech (TTS)
1 Introduction
Dysarthria is a neurological motor speech disorder primarily caused by damage to the central or peripheral nervous system that impairs control over the muscles used for speech production [1]. It often results in slurred, slow, or imprecise articulation that hinders communication and poses significant challenges for automatic speech recognition (ASR) and speech-based assistive technologies, due to the high variability of pathological speech and the limited availability of labeled data [2, 3]. Prior work primarily follows two main research directions.
The first direction focuses on ASR data augmentation, leveraging speech synthesis models to generate dysarthric speech and improve recognition robustness. The Rhythm-and-Voice (RnV) framework [4] employs self-supervised speech representations to disentangle and convert rhythmic and speaker-related characteristics in an unsupervised manner, and adapts ASR models using the converted speech. Soleymanpour et al. [5] propose a modified neural multi-talker TTS system that incorporates a dysarthria severity coefficient and a pause insertion module to synthesize pathological speech at varying severity levels for controlled data augmentation. Leung et al. [6] introduce a diffusion-based text-to-dysarthric-speech (TTDS) synthesis framework and fine-tune large-scale ASR models using the generated pathological data to improve recognition performance on dysarthric speech.
A second line of research focuses on dysarthric speech reconstruction, aiming to transform pathological speech into more intelligible speech. Early approaches adapt voice conversion (VC) techniques to map dysarthric speech to clearer acoustic realizations [7, 8]. E2E-DSR [9] replaces the text encoder in a sequence-to-sequence TTS system with a pre-trained ASR model, enabling end-to-end reconstruction from dysarthric speech to control-style speech. Unit-DSR [10] leverages self-supervised representations and discrete speech units to constrain content restoration within a linguistic space and improve training efficiency. More recently, Diff-DSR [11] adopts a latent diffusion model with content and speaker encoders to enhance reconstruction quality while preserving speaker identity.
Despite these advances, existing methods often entangle speaker identity with pathological articulation. This limits the diversity and controllability of synthesized speech for augmentation and can lead to speaker drift in reconstructed speech.
In this study, we propose ProtoDisent-TTS, a prototype-based disentanglement TTS framework built on a pre-trained Index-TTS model [12]. A pathology prototype codebook represents control and dysarthric articulation patterns, while a speaker encoder extracts timbre embeddings from a short prompt audio. These representations are combined with textual content for speech generation. A dual-classifier objective with a gradient reversal layer (GRL) [13] enforces the separation of speaker identity and pathology-related attributes within a unified generation pipeline.
The main contributions of this paper are summarized as:
  • We introduce ProtoDisent-TTS, a prototype-based TTS framework that enables controllable and bidirectional transformation between control and dysarthric speech within a unified latent space.
  • We demonstrate its effectiveness for scalable ASR data augmentation by generating pathology-controlled and speaker-consistent dysarthric speech.
  • We show that the same representation supports speaker-aware dysarthric speech reconstruction, bridging robust ASR training and assistive speech applications.
Figure 1: Overview architecture of ProtoDisent-TTS.
2 Methodology
2.1 Problem formulation and framework overview
We consider the problem of controllable dysarthric speech generation as follows. Given an input text sequence $x$, a prompt audio $a$ from a target speaker, and a speech condition label (prototype index) $k\in\{0,\dots,n\}$, where $k=0$ denotes healthy speech, $k\geq 1$ denotes one of several dysarthric prototypes, and $n$ is the total number of dysarthria prototypes, our goal is to synthesize a speech signal $\hat{y}$ that preserves the speaker’s vocal timbre and linguistic content while exhibiting articulation patterns specified by the selected pathology condition.
An overview of the proposed framework is illustrated in Figure 1 (a). A prompt audio $a$ is processed by the built-in encoder, comprising a conditioner and a perceiver module (hereafter referred to as the conditioner–perceiver), to extract a speaker embedding, $\mathbf{s}\in\mathbb{R}^{D},$ where $D$ denotes the dimensionality of the latent embedding space. In parallel, a pathology prototype embedding $\mathbf{p}_{k}\in\mathbb{R}^{D}$ is retrieved from a learnable codebook
$\mathcal{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{n}\},\quad\mathbf{p}_{i}\in\mathbb{R}^{D},$ (1)
according to the selected prototype index $k$.
During training, a dual-classifier is introduced to enforce a disentangled representation. The speaker embedding $\mathbf{s}$ is encouraged to encode speaker timbre while remaining invariant to dysarthria-related characteristics, whereas the pathology prototype embedding $\mathbf{p}_{k}$ is encouraged to capture the articulation patterns associated with the $k$-th dysarthria condition.
The speaker embedding $\mathbf{s}$ and the selected pathology prototype $\mathbf{p}_{k}$ are combined via element-wise addition to form a joint representation
$\mathbf{z}=\mathbf{s}+\mathbf{p}_{k},\quad\mathbf{z}\in\mathbb{R}^{D},$ (2)
which is then passed to the Index-TTS text-to-speech language model with text sequence $x$ for speech synthesis.
2.2 Prototype-based Representation and Disentanglement
To explicitly separate pathology-related articulation patterns from speaker identity, we introduce a prototype-based representation combined with a dual-classifier disentanglement objective. The pathology prototype codebook $\mathcal{P}$ consists of a set of learnable embeddings, where $\mathbf{p}_{0}$ represents a healthy (non-dysarthric) articulation prototype and $\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}$ correspond to distinct dysarthric articulation prototypes derived from patient groups. By conditioning generation on a selected prototype embedding, the framework provides an interpretable and controllable interface over pathological speech characteristics.
The conditioner–perceiver extracts a speaker embedding $\mathbf{s}$ from the prompt audio. Without explicit constraints, dysarthria-related information may leak into this representation. To prevent such entanglement, we employ a dual-classifier strategy that enforces complementary objectives on the speaker and prototype embeddings.
First, a dysarthria condition classifier $\mathcal{C}_{\text{dys}}$ operates on the combined representation $\mathbf{s}+\mathbf{p}_{k}$ and is trained using a binary classification objective that distinguishes between dysarthric and healthy speech conditions. This objective encourages pathology-related attributes to be captured primarily by the prototype embedding $\mathbf{p}_{k}$.
[…]
Together, the prototype codebook and the dual-classifier objective establish a factorized and interpretable latent space in which speaker identity and pathology-related attributes can be independently controlled and selectively substituted within a unified TTS framework. The overall training objective follows the standard Index-TTS formulation and minimizes a text-to-speech loss $L_{\text{TTS}}$, defined as the sum of cross-entropy losses over the predicted mel tokens and text tokens. In addition, cross-entropy losses are computed for both condition classifiers to enforce the proposed disentanglement constraints. The total training objective is defined as:
$L_{\text{Total}}=L_{\text{TTS}}+\alpha\,L_{\mathcal{C}_{\text{dys}}}+\beta\,L_{\mathcal{C}_{\text{adv}}},$ (3)
where $\alpha$ and $\beta$ weight the contributions of the dysarthria condition classifier and the adversarial classifier losses, respectively, relative to the text-to-speech loss.
Moreover, to further strengthen disentanglement, we incorporate a voice conversion–based training scheme [14] to construct cross-condition, timbre-shifted speech pairs. Specifically, dysarthric utterances are converted to the timbre of randomly selected healthy speakers, while healthy utterances are converted to the timbre of randomly selected dysarthric speakers. This design enables diverse combinations of articulation characteristics and speaker timbre, encouraging the model to learn representations in which speaker identity and pathology-related features are more clearly separated.
2.3 Prototype-based Control and Bidirectional Transformation
Owing to its prototype-based control mechanism, ProtoDisent-TTS enables the synthesis of speech by combining speaker timbre and dysarthric articulation patterns, making bidirectional transformation between healthy and dysarthric speech feasible. This capability provides a unified interface for both ASR data augmentation and dysarthric speech reconstruction.
For ASR augmentation, unlike prior approaches that primarily generate dysarthric speech from pathological speakers, ProtoDisent-TTS can synthesize dysarthric speech with the timbre of healthy speakers while imposing dysarthric characteristics defined by the selected pathology prototypes. In addition, the framework enables modification of the dysarthric characteristics of real pathological speech by substituting different prototypes, thereby generating diverse training samples that span multiple dysarthria patterns and severity levels. As illustrated in Figure 1(b), given a healthy speech sample and a randomly selected prototype index $k$, ProtoDisent-TTS synthesizes speech that preserves the healthy speaker’s timbre while exhibiting dysarthric articulation patterns specified by the selected prototype. The resulting synthetic dataset can then be used as an augmentation resource for ASR training.
For speech reconstruction, ProtoDisent-TTS transforms dysarthric speech into healthy, intelligible speech while preserving speaker identity through the retained speaker timbre embedding. As shown in Figure 1(c), a robust ASR model, enhanced through the proposed augmentation strategy, is first used to predict the textual content $\hat{x}$ from the input dysarthric speech. The recognized text $\hat{x}$ and the same dysarthric speech, used as prompt audio, are then passed to ProtoDisent-TTS to synthesize speaker-consistent, healthy speech.
3 Datasets and implementation details
3.1 Datasets
We use the TORGO database of dysarthric speech [15], which consists of aligned acoustic recordings and measured features from 8 English-speaking individuals (3 female and 5 male) diagnosed with either cerebral palsy (CP) or amyotrophic lateral sclerosis (ALS). In addition, a matched control group comprising 7 healthy speakers (3 female and 4 male) is included. The dysarthric speakers in the TORGO database were assessed by a speech-language therapist (SLT) using the Frenchay Dysarthria Assessment (FDA) [16]. Based on these clinical evaluations, the eight pathological speakers are categorized into four severity levels: severe, moderate–severe, moderate, and mild.
3.2 Implementation details
The proposed ProtoDisent-TTS is built upon a pre-trained Index-TTS backbone. We incorporate LoRA adapters [17] into the text-to-speech language model and enable parameter updates for the perceiver module (i.e., only the perceiver component of the conditioner–perceiver), the dual-classifier modules, and the pathology prototype codebook, while all remaining components of Index-TTS are kept frozen during fine-tuning.
For LoRA, we target the attention and MLP modules, and set the rank of the low-rank adaptation to $r=16$. The learning rate for the LoRA parameters is set to $5\times 10^{-5}$.
In addition, the pathology codebook, the dual-classifier modules, and the perceiver are jointly optimized during fine-tuning in separate parameter groups in the optimizer with learning rates of $2.5\times 10^{-3}$, $2.5\times 10^{-4}$, and $2.5\times 10^{-4}$, respectively. The weighting coefficients $\alpha$ and $\beta$ are both set to 1, and the gradient reversal layer (GRL) scaling factor $\lambda$ is set to 0.1. The total number of dysarthric prototypes $n$ in the pathology prototype codebook is set to the number of dysarthric patients plus one additional prototype that represents all healthy speakers, with each patient corresponding to a distinct prototype $k$.
4 Experiments
We conduct three targeted experiments to evaluate the effectiveness of ProtoDisent-TTS for ASR data augmentation and dysarthric speech reconstruction. Specifically, we evaluate: (1) Synthetic dysarthria data substitution, where pathological speech is synthesized using speaker-specific pathology prototypes to assess the feasibility of replacing real dysarthric data with fully synthetic training samples; (2) Healthy-to-dysarthric transformation for data augmentation, which combines healthy speaker timbre with diverse pathology prototypes to examine the impact of scalable synthetic data augmentation on ASR performance; and (3) Speaker identity preservation in dysarthric-to-healthy reconstruction, which evaluates whether speaker identity can be faithfully retained when reconstructing healthy speech from dysarthric inputs.
We adopt Whisper models [18] as the ASR backbone and employ a leave-one-speaker-out (LOSO) evaluation protocol [19, 6], in which the speech data from one speaker is the test set, while the data from remaining speakers is for training.
Figure 2: Comparison of ASR performance trained on real and synthetic dysarthric speech.
4.1 Synthetic dysarthria data substitution
This experiment evaluates whether synthetic dysarthric speech generated by ProtoDisent-TTS can serve as an effective substitute for real pathological speech in ASR training. We adopt Whisper-Medium and Whisper-Large as baseline ASR models and compare their performance under three conditions: (i) the pre-trained models used directly without task-specific adaptation (Whisper-Medium and Whisper-Large), (ii) fine-tuning using only real dysarthric speech data (Whisper-Medium-real and Whisper-Large-real), and (iii) fine-tuning using fully synthetic dysarthric speech generated by ProtoDisent-TTS in place of the real training set (Whisper-Medium-syn and Whisper-Large-syn).
The experiment is conducted on all patients with dysarthria, and performance is reported in terms of word error rate (WER) and phoneme error rate (PER). The quantitative results are presented in Fig. 2. While models trained on real data consistently achieve the best performance, training with synthetic data yields substantial improvements over the pre-trained models without adaptation, and the performance gap between real and synthetic training remains relatively small. This demonstrates the practical utility of the proposed synthetic data generation framework for building robust dysarthric ASR systems.
Table 1: Word error rate (WER) under different synthetic data augmentation ratios using Whisper-Medium (WM) on the severe dysarthria subset (F01, M01, M02, M04).
Training Setting F01 M01 M02 M04
WM + 0% 0.4161 0.3101 0.3866 0.5244
WM + 20% 0.3080 0.2152 0.2809 0.3682
WM + 40% 0.2866 0.1755 0.2705 0.3078
WM + 60% 0.2619 0.1873 0.2242 0.2719
WM + 80% 0.2553 0.1541 0.2500 0.2591
WM + 100% 0.2488 0.1535 0.2208 0.2631
Table 2: WER comparison with prior benchmarks on the TORGO ASR task.
Method FS2D FMLLR-DNN SD-CTL TTDS Ours
Severe. 0.5588 0.4329 0.6824 0.2330 0.2215
Mod-Severe 0.4960 0.4405 0.3315 0.1398 0.1174
Moderate 0.3680 0.3593 0.2284 0.0327 0.0419
Mild 0.1260 0.1165 0.1035 0.0257 0.0232
4.2 Healthy-to-dysarthric transformation for data augmentation
In the second experiment, we investigate how the proportion of synthetic data affects ASR performance when combined with real training data. Specifically, dysarthric speech is synthesized by combining healthy speaker prompts with different dysarthric prototypes. Starting from a fixed set of all real speech samples (dysarthria and healthy), we incrementally augment the training set with synthetic data at ratios of 20%, 40%, 60%, 80%, and 100%. Note that, we only focus on patient in severity group (F01, M01, M02, M04). Whisper-Medium models are trained under each configuration, and recognition performance is evaluated to analyze the scaling behavior and marginal benefits of synthetic data augmentation.
We report the WER results in Table 1. As the proportion of synthetic data increases, the WER consistently decreases, indicating improved recognition performance. A substantial performance gain is observed when introducing a small amount of synthetic data (20%) compared to the no-augmentation setting. Beyond 80% synthetic data, the performance improvements begin to saturate, with the best results achieved when the training set is fully augmented with synthetic data (100%).
Moreover, we use our strongest configuration—Whisper-Medium trained on all real data and augmented with 100% synthetic speech—to benchmark against prior methods on the TORGO ASR task, including FS2D [20], FMLLR-DNN [21], SD-CTL [22], and TTDS [6]. Evaluation is conducted across four severity subgroups: Severe (F01, M01, M02, M04), Mod-Severe (M05), Moderate (F03), and Mild (F04, M03). The mean WER results are reported in Table 2, where our method achieves the best performance across all groups except the Moderate group, in which TTDS attains a slightly lower WER.
Table 3: Speaker similarity (mean cosine similarity) between original dysarthric speech and reconstructed healthy speech.
Method F01 M01 M02 M04 M05
CMHR 0.1671 0.1598 0.1593 0.1803 0.1575
w/o Disent. 0.1519 0.1784 0.2197 0.1777 0.1703
Ours 0.3438 0.3667 0.3436 0.2965 0.3385
4.3 Speaker identity preservation in dysarthric-to-healthy reconstruction
The third experiment evaluates speaker identity preservation in dysarthric-to-healthy speech reconstruction. To isolate identity effects from ASR errors, we use ground-truth text as input and compute speaker similarity using cosine distance between ECAPA-TDNN embeddings [23] of the original dysarthric and reconstructed healthy speech.
As a reference, we introduce a Content-Matched Healthy Reference (CMHR) baseline, which measures similarity between dysarthric utterances and healthy utterances with identical textual content, providing a lower-bound estimate of identity preservation. We further include an ablation variant, w/o Disent., in which the dual-classifier objective is removed.
The results in Table 3 show that ProtoDisent-TTS consistently outperforms both baselines, demonstrating the effectiveness of the proposed disentanglement mechanism in preserving speaker identity during reconstruction.
5 Conclusion
This paper proposes ProtoDisent-TTS, a prototype-based disentanglement TTS framework built upon a pre-trained text-to-speech backbone that synthesizes realistic dysarthric speech. Through its explicit disentanglement design, the framework further enables bidirectional transformation between healthy and dysarthric speech.
Across all experimental settings, ProtoDisent-TTS consistently improves ASR performance while achieving strong speaker identity preservation in speech reconstruction. Given that dysarthric ASR is inherently data-limited, these results highlight the potential of controllable speech synthesis as a practical and scalable approach for enhancing both recognition robustness and assistive speech technologies.
Acknowledgments
This work was supported by The Hong Kong Polytechnic University Start-up Fund (Project ID: P0053210), The Hong Kong Polytechnic University Faculty Reserve Fund (Project ID: P0053738), an internal grant from The Hong Kong Polytechnic University (Project ID: P0048377), The Hong Kong Polytechnic University Departmental Collaborative Research Fund (Project ID: P0056428), The Hong Kong Polytechnic University Collaborative Research with World-leading Research Groups Fund (Project ID: P0058097) and Research Grants Council Collaborative Research Fund (Ref: C5033-24G).